### v1

# The asymptoticity of extremal length in Teichmüller space

Abstract.
We study the asymptotic behavior of extremal length along Teichmüller rays. Specifically, we determine the limit of extremal length along a Teichmüller ray and obtain an explicit expression for this limit, which complements a related formula established by Cormac Walsh. Building on this result and Kerckhoff’s formula, we establish a formula for the limiting Teichmüller distance between two points moving along arbitrary pairs of Teichmüller rays. Furthermore, we derive a necessary and sufficient condition for two Teichmüller rays to be asymptotic. Finally, by shifting the initial points of the Teichmüller rays along their associated Teichmüller geodesics, we show that the minimum of the limiting Teichmüller distance coincides with the detour metric between the endpoints of the rays on the horofunction boundary.
Key words and phrases:
Teichmüller space, extremal length, Teichmüller ray.
2020 Mathematics Subject Classification:
30F60, 32G15, 57M50.
The second author was partially supported by the National Natural Science Foundation of China through Grant 12271017.
1. Introduction
Extremal length is introduced by Ahlfors and Beurling [1] for families of curves in the plane. It is a conformal invariant and has the quasiconformal distortion property. Kerckhoff [10] studied the extremal lengths of simple closed curves on Riemann surfaces, thereby proving that the Thurston compactification and Teichmüller compactification are distinct. More significantly, he extended extremal length to the measured foliation space $\mathcal{MF}$ and showed a useful formula for the Teichmüller distance, known as Kerckhoff’s formula. Gardiner and Masur [5] further developed the extremal length geometry of Teichmüller space. They showed that the infimum of the product of the extremal lengths for two transversely measured foliations is realized by a unique Teichmüller geodesic. Furthermore, they gave a compactification of Teichmüller space by the extremal length in a manner analogous to Thurston compactification. Walsh [21] studied the asymptotic behavior of extremal length and established a formula for the limiting extremal length along a Teichmüller ray. In addition, he proved that the Gardiner-Masur compactification coincides with the Thurston compactification. This result was also established independently by Liu and Su [12].
Let $S$ be an oriented topological surface of genus $g$ with $m$ punctures such that $3g-3+m\geq 1$. The Teichmüller space $\mathcal{T}(S)$ of $S$ is the space of marked Riemann surfaces up to Teichmüller equivalence. It is endowed with a natural complete metric $d_{\mathcal{T}}(\cdot,\cdot)$, called the Teichmüller metric. The Teichmüller ray $\mathcal{R}_{q,X}(t)$ induced by a unit-norm quadratic differential $q$ on $X$ is a geodesic ray in $\mathcal{T}(S)$ with respect to the Teichmüller metric.
Let $V(q)$ and $H(q)$ denote the vertical and horizontal foliations of $q$ respectively. It is known that, by removing the critical graph $\Gamma_{q}$ of $V(q)$, the surface $X$ can be decomposed into finitely many cylinders swept out by closed leaves and minimal domains in which all leaves are dense. An indecomposable component of $V(q)$ is either a cylindrical component or a minimal component with a transverse ergodic measure. It can be shown that the number of projectively distinct ergodic measures on a minimal domain is finite.
Miyachi [16] showed that a Teichmüller ray admits a limit in the Gardiner-Masur compactification, and furthermore, that different Teichmüller rays from $X$ have different limits. In particular, Miyachi [15] provided an explicit formula for the limit in the case of Jenkins-Strebel rays, which is also proved by Kerckhoff [10]. In [21], Walsh extended this formula to arbitrary Teichmüller rays. For any measured foliation $F\in\mathcal{MF}$, Walsh showed that
$\lim_{t\to\infty}e^{-2t}\mathrm{Ext}_{\mathcal{R}_{q,X}(t)}(F)=\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F)^{2}}{i(G_{j},H(q))},$
where $G_{j}$ is an indecomposable component of $V(q)$ and $a_{j}$ is a positive coefficient. Furthermore, Walsh proved that the limit of a Teichmüller ray in the Gardiner-Masur boundary corresponds to a Busemann point in the horofunction boundary and vice versa. It was observed that non-Busemann points exist in the horofunction boundary, a result that was independently confirmed by Miyachi [17].
It is evident that the limit is degenerated if the measured foliation $F$ has no geometric intersection with $V(q)$. In order to ascertain the effect of the measured foliations that do not intersect with $V(q)$, it is necessary to consider the asymptotic behavior of $e^{2t}\mathrm{Ext}_{\mathcal{R}_{q,X}(t)}(\cdot)$ for such measured foliations. The explicit limit expression for these measured foliations are derived as follows.
Theorem 1.1.
Let $\mathcal{R}_{q,X}(t)$ be the Teichmüller ray from $X\in\mathcal{T}(S)$ determined by the unit-norm quadratic differential $q$. Suppose that the vertical foliation of $q$ can be expressed as $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}$ is a positive coefficient and $G_{j}$ is an indecomposable component of $V(q)$. Then for any $F\in\mathcal{MF}$,
$\lim_{t\to\infty}e^{2t}\mathrm{Ext}_{X_{t}}(F)=\sup_{F^{\prime}\in\mathcal{MF}}\frac{i(F,F^{\prime})^{2}}{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F^{\prime})^{2}}{i(G_{j},H(q))}},$
where $X_{t}$ is a point on the Teichmüller ray $\mathcal{R}_{q,X}(t)$.
The limit is infinite if the measured foliation $F$ intersects $V(q)$, since in this case the measured foliation $F^{\prime}$ can be chosen to coincide with $V(q)$. Moreover, if the measured foliation $F$ is similar to $V(q)$, that is, if $F$ can be expressed as $F=\sum_{j=1}^{n}c_{j}G_{j}$, where $c_{j}\geq 0$ is a coefficient, the right side of the formula can be simplified.
Corollary 1.2.
Let $\mathcal{R}_{q,X}(t)$ be the Teichmüller ray determined by the unit-norm quadratic differential $q$, and let $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ be the vertical foliation. If a measured foliation $F\in\mathcal{MF}$ can be written in the form $F=\sum_{j=1}^{n}c_{j}G_{j}$, where $c_{j}\geq 0$ is a coefficient, then
$\lim_{t\to\infty}e^{2t}\mathrm{Ext}_{X_{t}}(F)=\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q))}{a_{j}}.$
The long-term behavior of Teichmüller rays is a fundamental question in Teichmüller geometric. In this paper, we focus on the asymptotic behavior of two Teichmüller rays in Teichmüller space. The divergence case for two Teichmüller rays has been clarified by the efforts of Ivanov, Lenzhen and Masur. Ivanov [9] showed that two Teichmüller rays are divergent if their associated vertical foliations intersect. Furthermore, the Teichmüller rays remain bounded if the vertical foliations are absolutely continuous (see §2.2). The remaining cases were resolved by Lenzhen and Masur [11]. They proved that two Teichmüller rays are divergent if the vertical foliations are either topologically inequivalent or topologically equivalent but not absolutely continuous. Consequently, it can be concluded that two Teichmüller rays are bounded if and only if the vertical foliations are absolutely continuous.
For bounded Teichmüller rays, the remaining questions concern the limiting Teichmüller distances of pairs of Teichmüller rays and the conditions under which they are asymptotic. In [13], Masur considered the case of uniquely ergodic foliations, proving that two bounded Teichmüller rays with uniquely ergodic vertical foliations are asymptotic if their foliations contain no simple closed curves formed by saddle connections. Amano [3] derived a formula for the limiting Teichmüller distance of Jenkins-Strebel rays, showing that the limiting distance of bounded Jenkins-Strebel rays depends on both the moduli of cylindrical components of initial surfaces and the Teichmüller distance between their endpoints in the augmented Teichmüller spaces. In [7], we extended this formula to Teichmüller rays along which the surfaces can be decomposed into cylinders and minimal domains whose transverse measures are uniquely ergodic, with the decompsition determined by the leaves of the vertical foliations. By means of Kerckhoff’s formula, we determine the limiting distance for arbitrary pairs of Teichmüller rays through the asymptotic behavior of extremal lengths, thereby establishing an explicit expression for the limit for any such pair.
Theorem 1.3.
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by the unit-norm quadratic differentials $q$ and $q^{\prime}$ respectively.
  • (i)
    If the vertical foliations $V(q)$ and $V(q^{\prime})$ are absolutely continuous, i.e. $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$, where $a_{j},b_{j}>0$ are coefficients, then
    $\lim_{t\to\infty}d_{\mathcal{T}}(X_{t},Y_{t})=\frac{1}{2}\log\max_{1\leq j\leq n}\left\{\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))},\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right\}.$
  • (ii)
    Otherwise,
    $\lim_{t\to\infty}d_{\mathcal{T}}(X_{t},Y_{t})=+\infty.$
The notion of modulus for a cylinder extends naturally to any indecomposable component of $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$. We define the modulus of the indecomposable component $G_{j}$ of $V(q)$ by
$m_{j}=\frac{a_{j}}{i(G_{j},H(q))}.$
Given two absolutely continuous vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$, we say that $q$ and $q^{\prime}$ (or $V(q)$ and $V(q^{\prime})$) are modularly equivalent if
$\frac{a_{j}}{i(G_{j},H(q))}=C\frac{b_{j}}{i(G_{j},H(q^{\prime}))},$
for all $j$, where $C$ is a positive constant.
Corollary 1.4.
The Teichmüller rays $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ are asymptotic if and only if $q$ and $q^{\prime}$ are modularly equivalent.
The limiting Teichmüller distance between any pair of Teichmüller rays only depends on the ratios of the moduli of the indecomposable components at the initial points. Therefore, By appropriately shifting these initial points along their corresponding geodesics, the minimum of the limiting Teichmüller distance can be achieved, which is associated with the detour metric on the horofunction boundary of $\mathcal{T}(S)$.
Proposition 1.5.
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by the unit norm quadratic differentials $q$ and $q^{\prime}$ respectively. If the vertical foliations $V(q)$ and $V(q^{\prime})$ are absolutely continuous, then the minimum of the limiting Teichmüller distance is $\frac{1}{2}\delta(\mathcal{B}_{q},\mathcal{B}_{q^{\prime}})$, where $\delta$ is the detour metric and $\mathcal{B}_{q},\mathcal{B}_{q^{\prime}}$ are the Busemann points on the horofunction boundary of $\mathcal{T}(S)$ associated to $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ respectively.
The following corollary, which has been proved by Walsh in [21], is a direct combination of Corollary 1.4 and Proposition 1.5.
Corollary 1.6.
The Busemann points $\mathcal{B}_{q}$ and $\mathcal{B}_{q^{\prime}}$ are identical if and only if $q$ and $q^{\prime}$ are modularly equivalent.
This paper is organized as follows. In Section 2, we recall some background on Teichmüller spaces, measured foliations, quadratic differentials, extremal lengths and Teichmüller rays. In Section 3, we investigate the asymptotic behavior of extremal length along Teichmüller rays based on the results of Walsh [21]. Then we prove Theorem 1.1 and Corollary 1.2. In Section 4, the ratios of extremal lengths along pairs of Teichmüller rays is explored. Furthermore, by Kerckhoff’s formula, we prove Theorem 1.3 and Corollary 1.4. In Section 5, we review some background and results on the horofunction compactification of Teichmüller space and detour metric, and we prove Proposition 1.5 and Corollary 1.6.
2. Preliminaries
2.1. Teichmüller spaces
Let $X$ be a Riemann surface and $f$ be an orientation-preserving homeomorphism from $S$ onto $X$. The pair $(X,f)$ is called a marked Riemann surface, where $f$ serves as the marking of $X$. Two marked Riemann surfaces $(X_{1},f_{1})$ and $(X_{2},f_{2})$ are said to be Teichmüller equivalent if there exists a conformal homeomorphism $\phi:X_{1}\to X_{2}$ such that $f_{2}^{-1}\circ\phi\circ f_{1}$ is homotopic to the identity mapping on $S$. The Teichmüller equivalent classes containing $(X,f)$ is denoted by $[X,f]$, and the Teichmüller space $\mathcal{T}(S)$ of $S$ is defined as the collection of all such equivalence classes. For notational simplicity, we usually use the Riemann surface $X$ to represent the class $[X,f]$.
The Teichmüller distance between two Riemann surfaces $X_{1},X_{2}\in\mathcal{T}(S)$ is given by:
$d_{\mathcal{T}}(X_{1},X_{2})=\frac{1}{2}\inf_{h}\{\log K(h)\},$
where the infimum is over all quasiconformal mapping $h:X_{1}\to X_{2}$ such that $f_{2}^{-1}\circ h\circ f_{1}$ is homotopic to the identity mapping on $S$, and $K(h)$ denotes the maximal quasiconformal dilatation of $h$. The Teichmüller space $\mathcal{T}(S)$ with the Teichmüller metric $d_{\mathcal{T}}$ is a complete geodesic metric space.
2.2. Measured foliations
Let $P$ be a finite set composed of finite points and punctures on $S$. A measured foliation on $S$ consists of a finite set $P$ and an open covering $\{U_{j}\}$ of $S\setminus P$ with a real-valued $C^{1}$-function $v_{j}$ on each $U_{j}$ such that $|dv_{j}|=|dv_{i}|$ on $U_{j}\cap U_{i}$. Each $p\in P$ is called a singularity of the measured foliation, and there is a neighborhood $V_{p}$ of $p$ with complex coordinate $z$ such that $|dv_{j}|=|\operatorname{Im}(z^{k/2}dz)|$ on $U_{j}\cap V_{p}$ for some $k\in\mathbb{Z}$. If $p\in P$ is a puncture of $S$, $k\geq-1$. Otherwise, $k\geq 0$. We denote by $(\mathcal{F},\mu)$ the measured foliation on $S$, where $\mu$ is a transverse invariant measure which is given by $|dv_{j}|$ in $U_{j}$, and $\mathcal{F}$ denotes the underlying foliation structure without considering the measure. The leaves of $\mathcal{F}$ correspond to curves on $S$ along which the functions $v_{j}$ remain constant.
Let $\mathcal{S}$ be the set of homotopy classes of essential non-peripheral unoriented simple closed curves on $S$. The intersection number of a measured foliation $(\mathcal{F},\mu)$ and a $\alpha\in\mathcal{S}$ is defined by
$i((\mathcal{F},\mu),\alpha)=\inf_{\alpha^{\prime}\in\alpha}\int_{\alpha^{\prime}}d\mu,$
where the infimum is taken over all simple closed curves in $\alpha$. Two measured foliations $(\mathcal{F}_{1},\mu_{1})$ and $(\mathcal{F}_{2},\mu_{2})$ are equivalent if
$i((\mathcal{F}_{1},\mu_{1}),\alpha)=i((\mathcal{F}_{2},\mu_{2}),\alpha)$
holds for any $\alpha\in\mathcal{S}$. We denote by $F$ the equivalent class containing $(\mathcal{F},\mu)$. Therefore, the function $i(F,\cdot)$ is an element of $\mathbb{R}^{\mathcal{S}}_{\geq 0}$, where $\mathbb{R}^{\mathcal{S}}_{\geq 0}$ is the space of non-negative functions on $\mathcal{S}$ topologized by pointwise convergence. We denote by $\mathcal{MF}$ the space of equivalent classes of measured foliations on $S$ equipped with the weak topology induced by the intersection functions in $\mathbb{R}^{\mathcal{S}}_{\geq 0}$. The intersection function can extend continuously to $\mathcal{MF}\times\mathcal{MF}$ (cf. [4] and [18]).
A saddle connection of $F$ is a leaf of $F$ joining two singularities, which may be identical. The union of all saddle connections of $F$ is denoted by $\Gamma_{F}$. It is known that each connected component of $S\setminus\Gamma_{F}$ is either a cylinder foliated by closed leaves or a minimal domain in which each leaf is dense. For a minimal domain $\Omega$ of $F$, there exists finite number of ergodic transverse measures $\{\mu_{i}\}_{i=1}^{k}$ on $\Omega$, where $k$ is determined by the topology of $\Omega$, such that any transverse invariant measure $\mu$ on $\Omega$ can be expressed as $\mu=\sum_{i=1}^{k}a_{i}\mu_{i}$ for $a_{i}\geq 0$. The measured foliation $F$ is uniquely ergodic on a minimal domain $\Omega$ if the transverse measure on $\Omega$ is unique up to scalar multiplication. We say that a measured foliation $F$ is uniquely ergodic if $S\setminus\Gamma_{F}$ contains only one minimal domain with a uniquely ergodic transverse measure. Then, the measured foliation $F$ can be represented as
$F=\sum_{j=1}^{n}a_{j}G_{j},$
where $G_{j}$ is either a cylindrical component or a minimal domain endowed with an ergodic transverse measure. The coefficients satisfy $a_{j}>0$ when $G_{j}$ is a cylindrical component and $a_{j}\geq 0$ when $G_{j}$ is a minimal component.
Two measured foliations $F_{1},F_{2}\in\mathcal{MF}$ are topologically equivalent if there exists a homeomorphism $f:S\setminus\Gamma_{F_{1}}\to S\setminus\Gamma_{F_{2}}$ that is isotopic to the identity and maps each leaf of $F_{1}$ to a corresponding leaf of $F_{2}$. In this case, $F_{1}$ and $F_{2}$ can be expressed as
$F_{1}=\sum_{j=1}^{n}a_{j}G_{j},\quad F_{2}=\sum_{j=1}^{n}b_{j}G_{j},$
where $G_{j}$ is either a cylindrical component or a minimal component. Furthermore, we say that $F_{1}$ and $F_{2}$ are absolutely continuous if they are topologically equivalent and their corresponding coefficients with positive values coincide. Then, for absolutely continuous foliations, we retain only the components with positive coefficients such that $a_{j},b_{j}>0$ for all $j=1,\cdots,n$.
Since there is a natural bijection between measured foliations on any two surfaces in $\mathcal{T}(S)$, achieved through pullback of measured foliations to the corresponding measured foliations on $S$, we can extend the notions of topologically equivalent and absolutely continuous to pairs of measured foliations on distinct Riemann surfaces. For simplicity, we shall denote by $F\in\mathcal{MF}$ the corresponding measured foliation on any surface $X\in\mathcal{T}(S)$.
2.3. Quadratic differentials
A holomorphic quadratic differential $q$ on a Riemann surface $X$ is a tensor of the form $q(z)dz^{2}$, where $q$ is a holomorphic function that is allowed to have simple poles at the punctures of $X$. Each $q$ induces a singular flat metric $|q|=|q(z)||dz|^{2}$ on $X$. The zeros and poles of $q$ refer to the critical points of $q$, which form the singularities of the flat metric on $X$. The space of all holomorphic quadratic differentials on $X$ is denoted by $Q(X)$ and is a Banach space with the norm $\|q\|=\int_{X}|q(z)|dxdy$.
Let $q$ be a holomorphic quadratic differential on $X$ and $I\subset\mathbb{R}$ be an interval. A trajectory of $q$ is a smooth path $\gamma:I\to X$ on $X$ such that the $\arg{q(\gamma(t))\gamma^{\prime}(t)^{2}}$ is constant on $I$. A trajectory of $q$ is maximal if it is not a proper sub-trajectory of any trajectory on $X$. There exists a canonical coordinate chart $z=x+iy$ for each point in $X$ except for critical points of $q$, such that $q=dz^{2}$ in the chart. A vertical trajectory of $q$ is a maximal trajectory on $X$ such that it is a vertical line in each canonical coordinate chart. The collection of all critical trajectories of $q$ forms a measured foliation $V(q)$ on $X$, called the vertical foliation of $q$, whose transverse measure is given by $|dx|$. Similarly, a horizontal trajectory of $q$ is a maximal trajectory on $X$ which is a horizontal line in a canonical coordinate chart. The horizontal foliation $H(q)$ of $q$ is formed by all horizontal trajectories of $q$, equipped with the transverse measure induced by $|dy|$.
A critical trajectory of $V(q)$ is a vertical trajectory with an endpoint at a critical point of $q$. The critical graph $\Gamma_{q}$ of $V(q)$ is the union of all critical trajectories of $V(q)$ connecting critical points of $q$. A holomorphic quadratic differential $q$ is called a Jenkins-strebel differential if all the leaves of $V(q)$ are closed except for the critical graph $\Gamma_{q}$. Hubbard and Masur [8] showed that for each measured foliation $F\in\mathcal{MF}$, there exists a holomorphic quadratic differential $q$ on $X$ whose vertical foliation $V(q)$ realizes the measured foliation $F$.
2.4. Extremal length
We recall the extremal length of a simple closed curve $\alpha\in\mathcal{S}$ on $X,f]\in\mathcal{T}(S)$. Let $\rho=\rho(z)|dz|$ be a Borel measurable conformal metric on $X$. Under this metric $\rho$, the length of $\alpha\in\mathcal{S}$ on $X$ is
$\ell_{\rho}(\alpha)=\inf_{\gamma^{\prime}\in f(\alpha)}\int_{\gamma^{\prime}}\rho(z)|dz|,$
where $\gamma^{\prime}$ ranges over all simple closed curves in the homotopy class $f(\alpha)$ corresponding to $\alpha\in\mathcal{S}$. The area of $X$ is
$\mathrm{Area}_{\rho}(X)=\iint_{X}\rho(z)^{2}dxdy.$
The extremal length $\mathrm{Ext}_{X}(\alpha)$ of $\alpha$ on $X$ is defined by
$\mathrm{Ext}_{X}(\alpha)=\sup_{\rho}\frac{\ell_{\rho}(\alpha)^{2}}{\mathrm{Area}_{\rho}(X)},$
where $\rho$ takes over all Borel measurable conformal metric on $X$ with $\mathrm{Area}_{\rho}(X)<\infty$.
Kerckhoff [10] showed that there is a unique continuous extension of the extremal length function from $\mathcal{S}$ to $\mathcal{MF}$ satisfying
$\mathrm{Ext}_{X}(tF)=t^{2}\mathrm{Ext}_{X}(F),\quad t\in\mathbb{R}_{>0},$
for any $F\in\mathcal{MF}$ and $X\in\mathcal{T}(S)$. Moreover, Kerckhoff established a useful formula of the Teichmüller distance by extremal length.
Lemma 2.1 (Kerckhoff’s formula).
The Teichmüller distance between $X,Y\in\mathcal{T}(S)$ is
$d_{\mathcal{T}}(X,Y)=\frac{1}{2}\log\sup_{\alpha\in\mathcal{S}}\frac{\mathrm{Ext}_{Y}(\alpha)}{\mathrm{Ext}_{X}(\alpha)}.$
Remark.
Since the unique continuous extension of the extremal length function to $\mathcal{MF}$ and the density of the set $\{t\alpha\mid t\in\mathbb{R}_{>0},\alpha\in\mathcal{S}\}$ in $\mathcal{MF}$, the Teichmüller distance between $X,Y\in\mathcal{T}(S)$ is also given by
$d_{\mathcal{T}}(X,Y)=\frac{1}{2}\log\sup_{F\in\mathcal{MF}}\frac{\mathrm{Ext}_{Y}(F)}{\mathrm{Ext}_{X}(F)}.$
Let $\mathcal{PMF}$ denote the projective space of measured foliations and $[F]$ be the equivalent class in $\mathcal{PMF}$ containing $F\in\mathcal{MF}$. Then, we also have
$d_{\mathcal{T}}(X,Y)=\frac{1}{2}\log\sup_{[F]\in\mathcal{PMF}}\frac{\mathrm{Ext}_{Y}(F)}{\mathrm{Ext}_{X}(F)}.$
Extremal length and intersection number are continuous on $\mathcal{MF}$, and a useful inequality between them can be derived from the work by Gardiner and Masur in [5], which is also given by Minsky [14] and Ivanov [9].
Lemma 2.2.
Let $F$ and $G$ be two measured foliations in $\mathcal{MF}$ and $X\in\mathcal{T}(S)$. Then
$\mathrm{Ext}_{X}(F)\mathrm{Ext}_{X}(G)\geq i(F,G)^{2}.$
Moreover, the inequality will be an equality if and only if $F$ and $G$ are realized by the horizontal and vertical foliations of a holomorphic quadratic differential on $X$.
For any $F\in\mathcal{MF}$ and $X\in\mathcal{T}(S)$, there is a holomorphic quadratic differential $q$ on $X$ with $F$ as its vertical foliation by Hubbard and Masur [8]. Then, there exists a measure foliation $G$ which is the horizontal foliation of $q$, such that the equality in Lemma 2.2 holds for $G$.
2.5. Teichmüller rays
Consider the quasiconformal mappings between two Riemann surfaces $X,Y\in\mathcal{T}(S)$. Teichmüller’s theorem states that there exists a unique quasiconformal mapping that realizes the Teichmüller distance $d_{\mathcal{T}}(X,Y)$, known as the Teichmüller mapping. For the Teichmüller mapping $f:X\to Y$, its Beltrami coefficient $\mu_{f}$ has the form $\mu_{f}=\frac{K(f)-1}{K(f)+1}\frac{\bar{q}}{|q|}$, where $q$ is a unit norm holomorphic quadratic differential on $X$.
Let $f_{q,t}$ be the Teichmüller mapping from $X$ to $X_{t}$ determined by the unit norm holomorphic quadratic differential $q$ on $X$. There is a unit norm holomorphic quadratic differential $q_{t}$ on $X_{t}$, called the terminal differential. Then, in the canonical coordinate $z=x+iy$ of $q$ and the canonical coordinate of $q_{t}$, the Teichmüller mapping $f_{q,t}$ is given by
$z\mapsto e^{t}x+e^{-t}y,$
where $t=\frac{1}{2}\log K(f_{q,t})$ is the arc length parameter with respect to Teichmüller distance.
The Teichmüller geodesic $\mathcal{G}_{q,X}(t)$ through $X$ induced by a unit norm holomorphic quadratic differential $q$ can be defined by
[…]
where $f_{q,t}:X\to X_{t}$ is the Teichmüller mapping for the holomorphic quadratic differential $q$ on $X$. The Teichmüller ray emanating from $X$ and induced by $q$ is defined by $\mathcal{R}_{q,X}(t):=\left.\mathcal{G}_{q,X}\right|_{\mathbb{R}_{\geq 0}}(t)$.
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays.
  • (i)
    The two Teichmüller rays are divergent if
    $d_{\mathcal{T}}(X_{t},Y_{t})\to+\infty\text{ as }t\to\infty.$
  • (ii)
    The two Teichmüller rays are bounded if there exists a constant $M>0$ such that $d_{\mathcal{T}}(X_{t},Y_{t})<M$ for any $t\geq 0$.
  • (iii)
    The two Teichmüller rays are asymptotic if
    $\lim_{t\to\infty}\inf_{Y^{\prime}\in\mathcal{R}_{q^{\prime},Y}(t)}d_{\mathcal{T}}(X_{t},Y^{\prime})=0.$
    Then, there is a $\sigma\in\mathbb{R}$ such that $d_{\mathcal{T}}(X_{t},Y_{t+\sigma})\to 0$ as $t\to\infty$.
3. The asymptotic behavior of extremal length
We investigate the asymptotic behavior of $e^{2t}\mathrm{Ext}_{X_{t}}(F)$ for any $F\in\mathcal{MF}$ along a Teichmüller ray $\mathcal{R}_{q,X}(t)$. This study serves as a complement to the results concerning the asymptotic behavior of $e^{-2t}\mathrm{Ext}_{X_{t}}(F)$ in [21].
We adopt the convention that the supremum of a function is evaluated exclusively over the domain where the function is well-defined. Specifically, We consider $c/0=\infty$ to be well-defined for any $c>0$, while $0/0$ and $\infty/\infty$ remain undefined.
Lemma 3.1 ([21]).
Let $\mathcal{R}_{q,X}(t)$ be a Teichmüller ray induced by a unit norm holomorphic quadratic differential $q$ on $X$. The vertical foliation of $q$ is $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}\geq 0$. Then,
$e^{-2t}\mathrm{Ext}_{X_{t}}(F)\geq\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F)^{2}}{i(G_{j},H(q))},$
for any $F\in\mathcal{MF}$ and $t\geq 0$.
Theorem 3.2 ([21]).
Let $\mathcal{R}_{q,X}(t)$ be a Teichmüller ray induced by a unit norm holomorphic quadratic differential $q$ on $X$, and $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}\geq 0$. Then,
$\lim_{t\to\infty}e^{-2t}\mathrm{Ext}_{X_{t}}(F)=\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F)^{2}}{i(G_{j},H(q))},$
for any $F\in\mathcal{MF}$.
Based on the results given by Walsh in [21] and Lemma 2.2, we show the limit of $e^{2t}\mathrm{Ext}_{X_{t}}(F)$ for any $F\in\mathcal{MF}$ along a Teichmüller ray.
Proof of Theorem 1.1.
By Lemma 2.2, for any measured foliations $F,F^{\prime}\in\mathcal{MF}$, we have
$\mathrm{Ext}_{X_{t}}(F)\mathrm{Ext}_{X_{t}}(F^{\prime})\geq i(F,F^{\prime})^{2}.$
Then, by Theorem 3.2, there is
$\liminf_{t\to\infty}e^{2t}\mathrm{Ext}_{X_{t}}(F)\geq\liminf_{t\to\infty}\frac{i(F,F^{\prime})^{2}}{e^{-2t}\mathrm{Ext}_{X_{t}}(F^{\prime})}=\frac{i(F,F^{\prime})^{2}}{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F^{\prime})^{2}}{i(G_{j},H(q))}}.$
Therefore,
$\liminf_{t\to\infty}e^{2t}\mathrm{Ext}_{X_{t}}(F)\geq\sup_{F^{\prime}\in\mathcal{MF}}\frac{i(F,F^{\prime})^{2}}{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F^{\prime})^{2}}{i(G_{j},H(q))}}.$ (1)
For any $F\in\mathcal{MF}$ and $X_{t}$ on the Teichmüller ray $\mathcal{R}_{q,X}(t)$, there exists a $H_{t}\in\mathcal{MF}$ by Lemma 2.2, such that
$\mathrm{Ext}_{X_{t}}(F)\mathrm{Ext}_{X_{t}}(H_{t})=i(F,H_{t})^{2}.$
By Lemma 3.1, we have
$e^{2t}\mathrm{Ext}_{X_{t}}(F)=\frac{i(F,H_{t})^{2}}{e^{-2t}\mathrm{Ext}_{X_{t}}(H_{t})}\leq\frac{i(F,H_{t})^{2}}{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},H_{t})^{2}}{i(G_{j},H(q))}}\leq\sup_{F^{\prime}\in\mathcal{MF}}\frac{i(F,F^{\prime})^{2}}{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F^{\prime})^{2}}{i(G_{j},H(q))}}.$
Then,
$\limsup_{t\to\infty}e^{2t}\mathrm{Ext}_{X_{t}}(F)\leq\sup_{F^{\prime}\in\mathcal{MF}}\frac{i(F,F^{\prime})^{2}}{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F^{\prime})^{2}}{i(G_{j},H(q))}}.$ (2)
By combining (1) and (2), the existence of the limit of $e^{2t}\mathrm{Ext}_{X_{t}}(F)$ along the Teichmüller ray $\mathcal{R}_{q,X}(t)$ is obtained, and we have
$\lim_{t\to\infty}e^{2t}\mathrm{Ext}_{X_{t}}(F)=\sup_{F^{\prime}\in\mathcal{MF}}\frac{i(F,F^{\prime})^{2}}{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F^{\prime})^{2}}{i(G_{j},H(q))}}$
for any $F\in\mathcal{MF}$. ∎
Proof of Corollary 1.2.
If the measured foliation $F$ can be expressed as $F=\sum_{j=1}^{n}c_{j}G_{j}$, where $c_{j}\geq 0$, then, by the Cauchy-Schwarz inequality, there is
$\frac{i(F,F^{\prime})^{2}}{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F^{\prime})^{2}}{i(G_{j},H(q))}}=\frac{\left(\sum_{j=1}^{n}c_{j}i(G_{j},F^{\prime})\right)^{2}}{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F^{\prime})^{2}}{i(G_{j},H(q))}}\leq\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q))}{a_{j}}$
for any $F^{\prime}\in\mathcal{MF}$.
The equality holds if and only if there is a $F^{\prime}\in\mathcal{MF}$ such that
$a_{j}i(G_{j},F^{\prime})=c_{j}i(G_{j},H(q))$
for each indecomposable component $G_{j}$. This $F^{\prime}$ satisfying the condition can be obtained by appropriately adjusting the transverse measures of $H(q)$ in the intersection set with $G_{j}$, scaled by the factor $c_{j}/a_{j}$. Then,
$\sup_{F^{\prime}\in\mathcal{MF}}\frac{i(F,F^{\prime})^{2}}{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F^{\prime})^{2}}{i(G_{j},H(q))}}=\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q))}{a_{j}}.$
By Theorem 1.1, we have
$\lim_{t\to\infty}e^{2t}\mathrm{Ext}_{X_{t}}(F)=\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q))}{a_{j}}$
for any $F=\sum_{j=1}^{n}c_{j}G_{j}$ in $\mathcal{MF}$, where $c_{j}\geq 0$. ∎
4. The asymptotic behavior of pairs of Teichmüller rays
The divergence case of pairs of Teichmüller rays is clear by the works of Ivanov [9], Lenzhen and Masur [11]. Based on Kerckhoff’s formula, we study the limit of Teichmüller distance between two points moving along two bounded Teichmüller rays, respectively.
Given a measured foliation $F_{0}\in\mathcal{MF}$, the space of measured foliation space $\mathcal{MF}$ can be partitioned into two complementary subspaces based on the intersection properties of measured foliations in $\mathcal{MF}$ with respect to $F_{0}$. Then, we have
$\mathcal{MF}=\mathcal{MF}_{0}(F_{0})\cup\mathcal{MF}_{1}(F_{0}),$
where
$\mathcal{MF}_{0}(F_{0})=\{F\in\mathcal{MF}\mid i(F_{0},F)=0\},$
and
$\mathcal{MF}_{1}(F_{0})=\{F\in\mathcal{MF}\mid i(F_{0},F)\neq 0\}.$
If two measured foliations $F,F^{\prime}\in\mathcal{MF}$ are absolutely continuous, then it is evident that
$\mathcal{MF}_{0}(F)=\mathcal{MF}_{0}(F^{\prime})\text{ and }\mathcal{MF}_{1}(F)=\mathcal{MF}_{1}(F^{\prime}).$
Let $\mathcal{R}_{q,X}(t)$ be a Teichmüller ray determined by a unit norm holomorphic quadratic differential $q$ on $X$. The vertical foliation of $q$ is $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}\geq 0$. We define a function $\mathcal{E}_{q}:\mathcal{MF}\to\mathbb{R}_{\geq 0}$ associated with the Teichmüller ray, which is given by
$\mathcal{E}_{q}(F)=\left\{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F)^{2}}{i(G_{j},H(q))}\right\}^{\frac{1}{2}}$
for any $F\in\mathcal{MF}$.
Lemma 4.1.
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by unit norm quadratic differentials $q$ and $q^{\prime}$, respectively. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then
$\lim_{t\to\infty}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}=\frac{\mathcal{E}_{q^{\prime}}(F)^{2}}{\mathcal{E}_{q}(F)^{2}}\leq\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}$
for any $F\in\mathcal{MF}_{1}(V(q))$.
Proof.
By Theorem 3.2, for any $F\in\mathcal{MF}_{1}(V(q))$, there is
$\lim_{t\to\infty}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}=\lim_{t\to\infty}\frac{e^{-2t}\mathrm{Ext}_{Y_{t}}(F)}{e^{-2t}\mathrm{Ext}_{X_{t}}(F)}=\frac{\sum_{j=1}^{n}\frac{b_{j}i(G_{j},F)^{2}}{i(G_{j},H(q^{\prime}))}}{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F)^{2}}{i(G_{j},H(q))}}=\frac{\mathcal{E}_{q^{\prime}}(F)^{2}}{\mathcal{E}_{q}(F)^{2}}.$
By relabeling the indices, we suppose that for all $j=1,\cdots,n,$
$\frac{\frac{b_{j}}{i(G_{j},H(q^{\prime}))}}{\frac{a_{j}}{i(G_{j},H(q))}}\leq\frac{\frac{b_{1}}{i(G_{1},H(q^{\prime}))}}{\frac{a_{1}}{i(G_{1},H(q))}}.$
Therefore,
$\frac{a_{1}}{i(G_{1},H(q))}\sum_{j=1}^{n}\frac{b_{j}i(G_{j},F)^{2}}{i(G_{j},H(q^{\prime}))}\leq\frac{b_{1}}{i(G_{1},H(q^{\prime}))}\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F)^{2}}{i(G_{j},H(q))}.$
Then,
$\frac{\mathcal{E}_{q^{\prime}}(F)^{2}}{\mathcal{E}_{q}(F)^{2}}=\frac{\sum_{j=1}^{n}\frac{b_{j}i(G_{j},F)^{2}}{i(G_{j},H(q^{\prime}))}}{\sum_{j=1}^{n}\frac{a_{j}i(G_{j},F)^{2}}{i(G_{j},H(q))}}\leq\frac{\frac{b_{1}}{i(G_{1},H(q^{\prime}))}}{\frac{a_{1}}{i(G_{1},H(q))}}=\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}.$
Indeed, the inequality in Lemma 4.1 has been shown in [21]. Since this inequality is crucial for our discussion, we state it and provide a proof here. Furthermore, Walsh established the following equality.
Lemma 4.2 ([21]).
Let $q$ and $q^{\prime}$ be two unit norm holomorphic quadratic differentials on $X,Y\in\mathcal{T}(S)$, respectively. The vertical foliation of $q$ is $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$, where $a_{j}>0$. If the vertical foliation of $q^{\prime}$ can be expressed as $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$, where $b_{j}\geq 0$, then
$\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathcal{E}_{q^{\prime}}(F)^{2}}{\mathcal{E}_{q}(F)^{2}}=\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}.$
Otherwise, the supremum is $\infty$.
Theorem 4.3.
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by unit norm quadratic differentials $q$ and $q^{\prime}$, respectively. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then
$\lim_{t\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}=\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}.$
Proof.
By Theorem 3.2 and Lemma 4.2, we have
$\liminf_{t\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}$ $\geq\sup_{F\in\mathcal{MF}_{1}(V(q))}\liminf_{t\to\infty}\frac{e^{-2t}\mathrm{Ext}_{Y_{t}}(F)}{e^{-2t}\mathrm{Ext}_{X_{t}}(F)}$ (3)
$=\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathcal{E}_{q^{\prime}}(F)^{2}}{\mathcal{E}_{q}(F)^{2}}$
$=\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}.$
By contradiction, we assume that
$\limsup_{t\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}>\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}.$
For simplicity, let
$E=\limsup_{t\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}\text{ and }M=\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}.$
There exists a subsequence $\{t_{k}\}_{k=1}^{\infty}$ such that
$\lim_{k\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{Y_{t_{k}}}(F)}{\mathrm{Ext}_{X_{t_{k}}}(F)}=E>M+\frac{E-M}{3}.$
Then, there is a $N_{0}$ such that for any $k>N_{0}$, we have
$\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{Y_{t_{k}}}(F)}{\mathrm{Ext}_{X_{t_{k}}}(F)}>M+\frac{E-M}{3}.$
For any $k>N_{0}$, there exists a $F_{k}\in\mathcal{MF}_{1}(V(q))$ such that
$\frac{\mathrm{Ext}_{Y_{t_{k}}}(F_{k})}{\mathrm{Ext}_{X_{t_{k}}}(F_{k})}>\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{Y_{t_{k}}}(F)}{\mathrm{Ext}_{X_{t_{k}}}(F)}-\frac{E-M}{6}>M+\frac{E-M}{6}.$ (4)
By Lemma 4.1, for any $F\in\mathcal{MF}_{1}(V(q))$, we have
$\lim_{k\to\infty}\frac{\mathrm{Ext}_{Y_{t_{k}}}(F)}{\mathrm{Ext}_{X_{t_{k}}}(F)}\leq M.$
Then, for $\varepsilon=\frac{E-M}{6}$, there is a $N_{1}>N_{0}>0$ such that for any $k>N_{1}$,
$\frac{\mathrm{Ext}_{Y_{t_{k}}}(F)}{\mathrm{Ext}_{X_{t_{k}}}(F)}<M+\varepsilon=M+\frac{E-M}{6}.$
For each $k>N_{1}$, we choose the same $F_{k}\in\mathcal{MF}_{1}(V(q))$ as in (4). Therefore, we have
$\frac{\mathrm{Ext}_{Y_{t_{k}}}(F_{k})}{\mathrm{Ext}_{X_{t_{k}}}(F_{k})}<M+\frac{E-M}{6}$ (5)
for any $k>N_{1}$. We derive a contradiction between (4) and (5). Then, there is
$\limsup_{t\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}\leq\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}.$ (6)
By combining (3) and (6), we can obtain that
$\lim_{t\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}=\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}.$
Subsequently, we consider the ratio of extremal lengths along a pair of Teichmüller rays $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$, for measured foliations in $\mathcal{MF}_{0}(V(q))$.
Lemma 4.4.
Let $q$ and $q^{\prime}$ be two unit norm holomorphic quadratic differentials on $X,Y\in\mathcal{T}(S)$, respectively. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then
$\sup_{F\in\mathcal{MF}^{\prime}_{0}(V(q))}\frac{\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q^{\prime}))}{b_{j}}}{\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q))}{a_{j}}}=\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))},$
where $\mathcal{MF}^{\prime}_{0}(V(q))=\{F\in\mathcal{MF}\mid F=\sum_{j=1}^{n}c_{j}G_{j},\;c_{j}\geq 0\}$.
Proof.
By relabeling the indices, we can suppose that for any $j=1,\cdots,n$,
$\frac{\frac{i(G_{j},H(q^{\prime}))}{b_{j}}}{\frac{i(G_{j},H(q))}{a_{j}}}\leq\frac{\frac{i(G_{1},H(q^{\prime}))}{b_{1}}}{\frac{i(G_{1},H(q))}{a_{1}}}.$
Then, for any $F\in\mathcal{MF}^{\prime}_{0}(V(q))$, there is
$\frac{i(G_{1},H(q))}{a_{1}}\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q^{\prime}))}{b_{j}}\leq\frac{i(G_{1},H(q^{\prime}))}{b_{1}}\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q))}{a_{j}}.$
Therefore,
$\frac{\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q^{\prime}))}{b_{j}}}{\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q))}{a_{j}}}\leq\frac{\frac{i(G_{1},H(q^{\prime}))}{b_{1}}}{\frac{i(G_{1},H(q))}{a_{1}}}=\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}.$
Moreover, we can choose a measured foliation $F\in\mathcal{MF}^{\prime}_{0}(V(q))$ such that $c_{1}>0$ and $c_{j}=\varepsilon$ for all $j=2,\cdots,n$. Then, there is
$\frac{\frac{c_{1}^{2}i(G_{1},H(q^{\prime}))}{b_{1}}+\varepsilon^{2}\sum_{j=2}^{n}\frac{i(G_{j},H(q^{\prime}))}{b_{j}}}{\frac{c_{1}^{2}i(G_{1},H(q))}{a_{1}}+\varepsilon^{2}\sum_{j=2}^{n}\frac{i(G_{j},H(q))}{a_{j}}}\leq\frac{\frac{i(G_{1},H(q^{\prime}))}{b_{1}}}{\frac{i(G_{1},H(q))}{a_{1}}}=\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}.$
Since the $\varepsilon$ can be chosen arbitrarily small, we can conclude that
$\sup_{F\in\mathcal{MF}^{\prime}_{0}(V(q))}\frac{\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q^{\prime}))}{b_{j}}}{\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q))}{a_{j}}}=\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}.$
Theorem 4.5.
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays induced by unit norm quadratic differentials $q$ and $q^{\prime}$, respectively. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then
$\lim_{t\to\infty}\sup_{F\in\mathcal{MF}_{0}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}=\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}.$
Proof.
By Lemma 2.2, for any $F,F^{\prime}\in\mathcal{MF}$, there are
$\mathrm{Ext}_{X_{t}}(F)\mathrm{Ext}_{X_{t}}(F^{\prime})\geq i(F,F^{\prime})^{2}.$
Then, we have
$e^{2t}\mathrm{Ext}_{X_{t}}(F)\geq\frac{i(F,F^{\prime})^{2}}{e^{-2t}\mathrm{Ext}_{X_{t}}(F^{\prime})}.$
Moreover, for any $F\in\mathcal{MF}$, there exists a $H^{\prime}_{t}\in\mathcal{MF}$ such that
$\mathrm{Ext}_{Y_{t}}(F)\mathrm{Ext}_{Y_{t}}(H^{\prime}_{t})=i(F,H^{\prime}_{t})^{2}.$
Then, there is
$e^{2t}\mathrm{Ext}_{Y_{t}}(F)=\frac{i(F,H^{\prime}_{t})^{2}}{e^{-2t}\mathrm{Ext}_{Y_{t}}(H^{\prime}_{t})}.$
Therefore, we have
$\limsup_{t\to\infty}\sup_{F\in\mathcal{MF}_{0}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}$ $=\limsup_{t\to\infty}\sup_{F\in\mathcal{MF}_{0}(V(q))}\frac{e^{2t}\mathrm{Ext}_{Y_{t}}(F)}{e^{2t}\mathrm{Ext}_{X_{t}}(F)}$ (7)
$=\limsup_{t\to\infty}\sup_{F\in\mathcal{MF}_{0}(V(q))}\frac{\frac{i(F,H^{\prime}_{t})^{2}}{e^{-2t}\mathrm{Ext}_{Y_{t}}(H^{\prime}_{t})}}{e^{2t}\mathrm{Ext}_{X_{t}}(F)}$
$\leq\limsup_{t\to\infty}\sup_{F\in\mathcal{MF}_{0}(V(q))}\frac{\frac{i(F,H^{\prime}_{t})^{2}}{e^{-2t}\mathrm{Ext}_{Y_{t}}(H^{\prime}_{t})}}{\frac{i(F,H^{\prime}_{t})^{2}}{e^{-2t}\mathrm{Ext}_{X_{t}}(H^{\prime}_{t})}}$
$=\limsup_{t\to\infty}\frac{e^{-2t}\mathrm{Ext}_{X_{t}}(H^{\prime}_{t})}{e^{-2t}\mathrm{Ext}_{Y_{t}}(H^{\prime}_{t})}$
$\leq\limsup_{t\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{X_{t}}(F)}{\mathrm{Ext}_{Y_{t}}(F)}.$
By contradiction, we assume that
$\limsup_{t\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{X_{t}}(F)}{\mathrm{Ext}_{Y_{t}}(F)}>\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}.$
For simplicity, let
$E^{\prime}=\limsup_{t\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{X_{t}}(F)}{\mathrm{Ext}_{Y_{t}}(F)}\text{ and }M^{\prime}=\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}.$
There exists a subsequence $\{t_{k}\}_{k=1}^{\infty}$ such that
$\lim_{k\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{X_{t_{k}}}(F)}{\mathrm{Ext}_{Y_{t_{k}}}(F)}=E^{\prime}>M^{\prime}+\frac{E^{\prime}-M^{\prime}}{3}.$
Then, there is a $N^{\prime}_{0}$ such that for any $k>N^{\prime}_{0}$, we have
$\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{X_{t_{k}}}(F)}{\mathrm{Ext}_{Y_{t_{k}}}(F)}>M^{\prime}+\frac{E^{\prime}-M^{\prime}}{3}.$
For any $k>N^{\prime}_{0}$, there exists a $F_{k}\in\mathcal{MF}_{1}(V(q))$ such that
$\frac{\mathrm{Ext}_{X_{t_{k}}}(F_{k})}{\mathrm{Ext}_{Y_{t_{k}}}(F_{k})}>\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{X_{t_{k}}}(F)}{\mathrm{Ext}_{Y_{t_{k}}}(F)}-\frac{E^{\prime}-M^{\prime}}{6}>M^{\prime}+\frac{E^{\prime}-M^{\prime}}{6}.$ (8)
By Lemma 4.1, for any $F\in\mathcal{MF}_{1}(V(q))$, there is
$\lim_{t\to\infty}\frac{\mathrm{Ext}_{X_{t}}(F)}{\mathrm{Ext}_{Y_{t}}(F)}\leq M^{\prime}.$
Then, for the subsequence $\{t_{k}\}_{k=1}^{\infty}$, we still have
$\lim_{k\to\infty}\frac{\mathrm{Ext}_{X_{t_{k}}}(F)}{\mathrm{Ext}_{Y_{t_{k}}}(F)}\leq M^{\prime}.$
Then, for $\varepsilon=\frac{E^{\prime}-M^{\prime}}{6}$, there is a $N^{\prime}_{1}>N^{\prime}_{0}>0$ such that for any $k>N^{\prime}_{1}$,
$\frac{\mathrm{Ext}_{X_{t_{k}}}(F)}{\mathrm{Ext}_{Y_{t_{k}}}(F)}<M^{\prime}+\varepsilon=M^{\prime}+\frac{E^{\prime}-M^{\prime}}{6}.$
For each $k>N^{\prime}_{1}$, we choose the same $F_{k}\in\mathcal{MF}_{1}(V(q))$ as in (8). Therefore, we have
$\frac{\mathrm{Ext}_{X_{t_{k}}}(F_{k})}{\mathrm{Ext}_{Y_{t_{k}}}(F_{k})}<M^{\prime}+\frac{E^{\prime}-M^{\prime}}{6}$ (9)
for any $k>N^{\prime}_{1}$. We derive a contradiction between (8) and (9). Then, there is
$\limsup_{t\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{X_{t}}(F)}{\mathrm{Ext}_{Y_{t}}(F)}\leq\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}.$ (10)
By combining (7) and (10), we can obtain that
$\limsup_{t\to\infty}\sup_{F\in\mathcal{MF}_{0}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}\leq\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}.$ (11)
Moreover, by Corollary 1.2 and Lemma 4.4, we have
$\liminf_{t\to\infty}\sup_{F\in\mathcal{MF}_{0}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}$ $\geq\sup_{F\in\mathcal{MF}_{0}(V(q))}\liminf_{t\to\infty}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}$ (12)
$\geq\sup_{F\in\mathcal{MF}^{\prime}_{0}(V(q))}\liminf_{t\to\infty}\frac{e^{2t}\mathrm{Ext}_{Y_{t}}(F)}{e^{2t}\mathrm{Ext}_{X_{t}}(F)}$
$=\sup_{F\in\mathcal{MF}^{\prime}_{0}(V(q))}\frac{\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q^{\prime}))}{b_{j}}}{\sum_{j=1}^{n}\frac{c_{j}^{2}i(G_{j},H(q))}{a_{j}}}$
$=\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}.$
By combining (11) and (12), we conclude that
$\lim_{t\to\infty}\sup_{F\in\mathcal{MF}_{0}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}=\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}.$
Proof of Theorem 1.3.
If the vertical foliations $V(q)$ and $V(q^{\prime})$ are absolutely continuous, then by Theorem 4.3, Theorem 4.5 and Kerckhoff’s formula, we have
$\lim_{t\to\infty}d_{\mathcal{T}}(X_{t},Y_{t})$ $=\lim_{t\to\infty}\frac{1}{2}\log\sup_{F\in\mathcal{MF}}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}$
$=\frac{1}{2}\log\lim_{t\to\infty}\max\left\{\sup_{F\in\mathcal{MF}_{0}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)},\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}\right\}$
$=\frac{1}{2}\log\max\left\{\lim_{t\to\infty}\sup_{F\in\mathcal{MF}_{0}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)},\lim_{t\to\infty}\sup_{F\in\mathcal{MF}_{1}(V(q))}\frac{\mathrm{Ext}_{Y_{t}}(F)}{\mathrm{Ext}_{X_{t}}(F)}\right\}$
$=\frac{1}{2}\log\max\left\{\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))},\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right\}$
$=\frac{1}{2}\log\max_{1\leq j\leq n}\left\{\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))},\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right\}.$
Otherwise, by the results of Ivanov [9], Lenzhen and Masur [11], the Teichmüller distance $d_{\mathcal{T}}(X_{t},Y_{t})$ tends to infinity as $t\to\infty$. ∎
Proof of Corollary 1.4.
If the Teichmüller rays $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ are asymptotic, then without loss of generality, we assume that their initial points are appropriately selected such that
$\lim_{t\to\infty}d_{\mathcal{T}}(X_{t},Y_{t})=0.$
Therefore, by Theorem 1.3, we have
$\frac{a_{j}}{i(G_{j},H(q))}=\frac{b_{j}}{i(G_{j},H(q^{\prime}))},$
for all $j=1,\cdots,n$. Then $q$ and $q^{\prime}$ are modularly equivalent.
Conversely, if $q$ and $q^{\prime}$ are modularly equivalent, then there exists a constant $C>0$ such that
$\frac{a_{j}}{i(G_{j},H(q))}=C\frac{b_{j}}{i(G_{j},H(q^{\prime}))},$
for all $j=1,\cdots,n$. Let $\sigma=\frac{1}{2}\log C$. There is
$\lim_{t\to\infty}d_{\mathcal{T}}(X_{t},Y_{t+\sigma})=\frac{1}{2}\log\max_{1\leq j\leq n}\left\{\frac{a_{j}i(G_{j},H(q^{\prime}))}{e^{2\sigma}b_{j}i(G_{j},H(q))},\frac{e^{2\sigma}b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right\}=0.$
Then the Teichmüller rays $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ are asymptotic. ∎
5. The minimum of the limiting Teichmüller distance
Since the limiting Teichmüller distance between two points along a pair of Teichmüller rays depends on the initial points of these rays, the minimum of the limiting Teichmüller distance can be attained by shifting the initial points along the corresponding Teichmüller geodesics. This minimum value is related to the detour metric $\delta$ between the two endpoints of the Teichmüller rays on the horofunction boundary of $\mathcal{T}(S)$.
5.1. The horofunction compactification
We recall the horofunction compactification of a metric space which is introduced by Gromov in [6]. Consider a proper geodesic metric space $(M,d)$ and choose a basepoint $b\in M$. A metric space is proper if all of its closed balls are compact, and it is geodesic if any two points can be connected by a geodesic segment. Define a function $\psi_{z}:M\to\mathbb{R}$ which is given by
$\psi_{z}(x):=d(x,z)-d(b,z),\;\forall x\in M.$
Then, this defines a mapping
$\begin{array}[]{cccc}\Psi:&M&\to&C(M)\\ &z&\mapsto&\psi_{z},\end{array}$
where $C(M)$ denotes the space of all continuous functions on $M$, equipped with the topology of uniform convergence on any compact subset of M.
Gromov showed that the mapping $\Psi$ is an embedding and the set $\Psi(M)$ is relatively compact in $C(M)$. The closure $\overline{\Psi(M)}$ is called the horofunction compactification of $M$. The horofunction boundary of $M$ is defined as $\partial_{hor}M:=\overline{\Psi(M)}\setminus\Psi(M)$, and each $\xi\in\partial_{hor}M$ is called a horofunction.
The horofunction boundary of $M$ depends on the choice of basepoint in $M$, and the horofunction boundaries of $M$ from different basepoints are homeomorphic. This homeomorphism is realized by applying additive constant adjustment to each horofunction.
Since the Teichmüller space $\mathcal{T}(S)$ with Teichmüller metric $d_{\mathcal{T}}$ is known to be a proper geodesic metric space, we can consider its horofunction compactification, denoted by $\overline{\mathcal{T}(S)}^{hor}$. The horofunction boundary of $\mathcal{T}(S)$ is denoted by $\partial_{hor}\mathcal{T}(S)$.
The Gardiner-Masur compactification of Teichmüller space is introduced by Gardiner and Masur in [5]. By using extremal lengths of simple closed curves, we can define a mapping
$\begin{array}[]{cccl}\varphi:&\mathcal{T}(S)&\to&\mathbb{R}_{\geq 0}^{\mathcal{S}}\\ &X&\mapsto&\left\{\mathrm{Ext}_{X}(\alpha)^{\frac{1}{2}}\right\}_{\alpha\in\mathcal{S}}.\end{array}$
Gardiner and Masur proved that the mapping $\Phi:=\pi\circ\phi:\mathcal{T}(S)\to P\mathbb{R}_{\geq 0}^{\mathcal{S}}$ is an embedding, where $\pi:\mathbb{R}_{\geq 0}^{\mathcal{S}}\setminus\{0\}\to P\mathbb{R}_{\geq 0}^{\mathcal{S}}$ is the natural projection, and the closure $\overline{\Phi(\mathcal{T}(S))}$ is compact. The Gardiner-Masur compactification of $\mathcal{T}(S)$, denoted by $\overline{\mathcal{T}(S)}^{GM}$, is defined by the closure $\overline{\Phi(\mathcal{T}(S))}$, and the Gardiner-Masur boundary of $\mathcal{T}(S)$ is defined as $\partial_{GM}\mathcal{T}(S):=\overline{\Phi(\mathcal{T}(S))}\setminus\Phi(\mathcal{T}(S))$.
Liu and Su [12] showed that the horofunction compactification of Teichmüller space is homeomorphic to the Gardiner-Masur compactification, which was also proved independently by Walsh in [21]. Walsh also showed that the Teichmüller ray $\mathcal{R}_{q,X}(t)$ converges to the projective class of $\mathcal{E}_{q}(\cdot)$ in $\overline{\mathcal{T}(S)}^{GM}$.
5.2. The detour metric
Let $\mathcal{I}$ be an unbounded subset of $\mathbb{R}_{\geq 0}$ containing 0. An almost-geodesic ray in a metric space $(M,d)$ is defined as a mapping $\gamma:\mathcal{I}\to M$ satisfying that for any $\varepsilon>0$, there exists a $T\geq 0$ such that
$|d(\gamma(0),\gamma(s))+d(\gamma(s),\gamma(t))-t|<\varepsilon,$
for all $s,t\in E$ with $t\geq s\geq T$.
Rieffel [19] proved that each almost-geodesic ray in a metric space $(M,d)$ converges to a point in $\partial_{hor}M$. A horofunction in $\partial_{hor}M$ is called a Busemann point if it is the limit of an almost-geodesic ray. The set of all Busemann points in $\partial_{hor}M$, called the Busemann set of $M$, is denoted by $\partial_{B}M$.
For any two horofunctions $\xi,\eta\in\partial_{hor}M$, the detour cost is defined by
$H(\xi,\eta):=\sup_{W\ni\xi}\inf_{x\in W}\left(d(b,x)+\eta(x)\right),$
where $W$ ranges over all neighborhoods of $\xi$ in the horofunction compactification of $(M,d)$. An equivalent definition is given by
$H(\xi,\eta):=\inf_{\gamma}\liminf_{t\to\infty}\left(d(b,\gamma(t))+\eta(\gamma(t))\right),$
where the infimum is taken over all paths $\gamma:\mathbb{R}_{\geq 0}\to M$ converging to $\xi$.
For any two Busemann points $\xi,\eta\in\partial_{B}M$, the detour distance between them is defined as
$\delta(\xi,\eta)=H(\xi,\eta)+H(\eta,\xi).$
We refer to [20] for more details. The detour metric $\delta$ may take infinity, and it is independent of the basepoint in $M$.
Walsh [21] showed that for any $X\in\mathcal{T}(S)$ and $\xi\in\partial_{B}\mathcal{T}(S)$, there exists a unique Teichmüller ray starting from $X$ and converging to $\xi$. We denote by $\mathcal{B}_{q}$ the Busemann point in $\partial_{B}\mathcal{T}(S)$, which is the limit of a Teichmüller ray $\mathcal{R}_{q,X}(t)$. Walsh also established a formula of detour distance between two Busemann points which are the limits of two Teichmüller rays from the basepoint of $\overline{\mathcal{T}(S)}^{hor}$ (see the Corollary $3$ in [21]). Amano [2] generalized this formula to any two Busemann points associated with distinct Teichmüller rays originating from different satrting points.
Proposition 5.1 ([2]).
Let $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ be two Teichmüller rays. If the vertical foliations $V(q)=\sum_{j=1}^{n}a_{j}G_{j}$ and $V(q^{\prime})=\sum_{j=1}^{n}b_{j}G_{j}$ are absolutely continuous, where $a_{j},b_{j}>0$, then the detour distance between $\mathcal{B}_{q}$ and $\mathcal{B}_{q^{\prime}}$ associated with these two rays is
$\delta(\mathcal{B}_{q},\mathcal{B}_{q^{\prime}})=\frac{1}{2}\log\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}+\frac{1}{2}\log\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}.$
If $V(q)$ and $V(q^{\prime})$ are not absolutely continuous, then $\delta(\mathcal{B}_{q},\mathcal{B}_{q^{\prime}})=+\infty$.
Proof of Proposition 1.5.
If $V(q)$ and $V(q^{\prime})$ are absolutely continuous, then by Theorem 1.3 and Proposition 5.1, we have
$\lim_{t\to\infty}d_{\mathcal{T}}(X_{t},Y_{t})$ $=\frac{1}{2}\log\max_{1\leq j\leq n}\left\{\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))},\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right\}$
$\geq\frac{1}{2}\log\left(\left(\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}\right)^{\frac{1}{2}}\cdot\left(\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right)^{\frac{1}{2}}\right)$
$=\frac{1}{2}\left(\frac{1}{2}\log\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}+\frac{1}{2}\log\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right)$
$=\frac{1}{2}\delta(\mathcal{B}_{q},\mathcal{B}_{q^{\prime}}).$
Since the detour distance $\delta(\mathcal{B}_{q},\mathcal{B}_{q^{\prime}})$ is invariant when the initial points of the Teichmüller rays $\mathcal{R}_{q,X}(t)$ and $\mathcal{R}_{q^{\prime},Y}(t)$ are shifted along their corresponding Teichmüller geodesics, then for any $\sigma\in\mathbb{R}$, there is
$\lim_{t\to\infty}d_{\mathcal{T}}(X_{t},Y_{t+\sigma})\geq\frac{1}{2}\delta(\mathcal{B}_{q},\mathcal{B}_{q^{\prime}}),$
where the equality holds if
$\sigma=\frac{1}{4}\log\frac{\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}}{\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}}.$
Thus, for this $\sigma$, we can obtain that
$\max_{1\leq j\leq n}\frac{e^{2\sigma}b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}$ $=\frac{\left(\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}\right)^{\frac{1}{2}}}{\left(\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right)^{\frac{1}{2}}}\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}$
$=\left(\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}\right)^{\frac{1}{2}}\cdot\left(\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right)^{\frac{1}{2}}$
$=\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{e^{2\sigma}b_{j}i(G_{j},H(q))}.$
Therefore, we conclude that
$\lim_{t\to\infty}d_{\mathcal{T}}(X_{t},Y_{t+\sigma})$ $=\frac{1}{2}\log\max_{1\leq j\leq n}\left\{\frac{a_{j}i(G_{j},H(q^{\prime}))}{e^{2\sigma}b_{j}i(G_{j},H(q))},\frac{e^{2\sigma}b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right\}$
$=\frac{1}{2}\log\left(\left(\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}\right)^{\frac{1}{2}}\cdot\left(\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right)^{\frac{1}{2}}\right)$
$=\frac{1}{2}\left(\frac{1}{2}\log\max_{1\leq j\leq n}\frac{a_{j}i(G_{j},H(q^{\prime}))}{b_{j}i(G_{j},H(q))}+\frac{1}{2}\log\max_{1\leq j\leq n}\frac{b_{j}i(G_{j},H(q))}{a_{j}i(G_{j},H(q^{\prime}))}\right)$
$=\frac{1}{2}\delta(\mathcal{B}_{q},\mathcal{B}_{q^{\prime}}).$
Proof of Corollary 1.6.
By Proposition 1.5, the Busemann points $\mathcal{B}_{q}$ and $\mathcal{B}_{q^{\prime}}$ are identical if and only if their associated Teichmüller rays are asymptotic. Combining this with Corollary 1.4, we can obtain that the Busemann points $\mathcal{B}_{q}$ and $\mathcal{B}_{q^{\prime}}$ are identical if and only if $q$ and $q^{\prime}$ are modularly equivalent. ∎
Acknowledgments
We are very grateful to Hideki Miyachi and Guangming Hu for their helpful discussions and invaluable encouragements.